OpenStudy (anonymous):

35 kilometers to centimeters? I need help please

OpenStudy (anonymous):

that's easy. how many centimeters in meter? How does to compare to how many kilometers there are in a meter?

OpenStudy (anonymous):

I don't know, I don't have no reference sheets

OpenStudy (anonymous):

well there are 1*10^2=100 centimeters in a meter and 1*10^-3 kilometers in a meter. this means that kilometers are larger than meters.

OpenStudy (anonymous):

considering meters is in between:centimeters<meters<kilometers, how many decimal places do you think we have to move, and in which direction?

OpenStudy (anonymous):

1km=10^5cm with means 35km=3500000cm

OpenStudy (anonymous):

Yes that's correct! just practice these kind of conversions and soon you'll master all of them. I suggest printing out a table with all the power as such: